Monday, January 31, 2011

Nightlog 2011-01-24

SA: Paul
SO: Patrick
Others: Siphelo, Zolisa, Darragh, Lisa, Charl, Hitesh


- The SAC baffle was reinstalled today. From the out of focus images we seemed
to see multiple spider components which we thought was due to the SAC baffle.
Took the SAC baffle off, but the multiple component remained. This cannot be
explained. The SAC baffled seemed well enough aligned in the way it was
installed. Multi component spider needs thinking about. The SAC baffle will be
re-installed tomorrow during the day.

- timdimm ran for a while but it was plagued by tracking issues.

- we continued the training and focused on getting the mirror alignment times
down to around 15min.

- some pointing offset observations were done and the results from that looks


No comments:

Post a Comment